Main page‎ > ‎Observing Runs‎ > ‎Night Reports‎ > ‎

Night report 2017-05-17

Night report 2017-05-17

Observers: Alessandro Razza, Felipe Olivares.
TAT on duty: Michel Dennefeld, Cosimo Inserra, Giorgos Leloudas.
Support team: Lluis Galbany.

Daytime calibrations: PESSTO_calib, PESSTO_bias
Owing to bad weather conditions skyflats were not obtained.

Weather conditions
Thick fog, Humidity > 90%, low atm. pressure (~ 760 hPa) the whole night.

Dome closed the whole night for the weather conditions.
Comments